OpenStudy (anonymous):

What is the product of the polynomials below? (6x3 - 3x2 + 5x + 9)(7x + 9)

OpenStudy (anonymous):

= 7x ( 6x3 - 3x2 + 5x + 9) + 9 (6x3 - 3x2 + 5x + 9)

OpenStudy (anonymous):

i dont ave no calculator around me

OpenStudy (anonymous):

No need calculator with simple multiply 1 by 1 digit, any way!

OpenStudy (anonymous):

42x4 + 33x3 + 8x2 + 108x + 81

OpenStudy (anonymous):

. 42x4 + 33x3 + 8x2 + 18x + 81

OpenStudy (anonymous):

The first term is correct! Do the same for all other terms!

OpenStudy (anonymous):

42x4 + 33x3 + 8x2 + 108x - 81

OpenStudy (anonymous):

42x4 + 33x3 + 8x2 + 108x + 81 i think its this

OpenStudy (anonymous):

Only your first term is correct, You can use calculator from you PC!

OpenStudy (anonymous):

7 * 3 - 7 * 5 = 7 * 9 =

OpenStudy (anonymous):

A. 42x4 + 33x3 + 8x2 + 18x + 81 B. 42x4 + 33x3 + 8x2 + 108x - 81 C. 42x4 + 33x3 + 8x2 + 108x + 81 D. 42x4 - 33x3 - 8x2 + 108x + 81 A. 42x4 + 33x3 + 8x2 + 18x + 81 these are the choices B. 42x4 + 33x3 + 8x2 + 108x - 81 C. 42x4 + 33x3 + 8x2 + 108x + 81 D. 42x4 - 33x3 - 8x2 + 108x + 81